### v1

@@ 5 Recommendations
Expanding the Scope of Disaster Management Features in Mobile Applications: In disaster management applications, the focus often lies on functionalities that are active during or immediately after a disaster, such as real-time alerts, disaster maps, and updates. While these are undoubtedly crucial, there is a significant opportunity to expand the feature set to better prepare and support users throughout the entire disaster lifecycle. For example, early warning alerts, which are present in only 22% of the surveyed applications, can provide critical advance notice to users, allowing them to take necessary precautions and potentially save lives.
@@
Moreover, the recovery and mitigation phases are equally important, yet often underrepresented in current applications. Features like shelter information, safety confirmations, and recovery actions are vital for helping individuals locate assistance, confirm the safety of friends and family, and understand the steps they can take next. Enhancing the availability of these features could significantly improve the efficacy of disaster response and recovery efforts, addressing gaps that currently leave users under prepared or insufficiently supported after a disaster strikes. By integrating a more comprehensive set of features that span the pre and post-disaster phases, developers can create more robust applications that cater to a wider range of user needs and scenarios.
@@
Internationalisation: As shown in Table 3, among the various apps categorized based on their geographic coverage—19 international apps, 6 region-specific apps, 19 national apps, and 6 territorial apps—only the international, region-specific, and territorial apps consistently support the English language. Among the national apps from non-English speaking countries, such as ‘Anhaar’ from Mongolia, ‘Emergency Ready’ App from South Korea, ‘BD 999’ from Bangladesh, ‘Alertswiss’ from Switzerland, ‘NINA - Die Warn-App des BBK’ from Germany, and several from Japan including ‘NERV Disaster Prevention’, ‘Safety Tips’, ‘National Evacuation Center Guide’, and ‘Yurekuru Call’, only ‘Safety Tips’, ‘NERV Disaster Prevention’, and ‘Emergency Ready’ App offer English language support.
@@
These examples underline the importance of incorporating at least one universally understood language, like English, in disaster management apps. This feature is essential for ensuring the usability of the app especially for visitors and temporary residents. Providing quality translations or, alternatively, enabling features that assist with language translation, like text copying, can significantly enhance the functionality and user experience of these apps.
@@
Addressing Topics with the Highest Percentage of Negative Reviews: From Section 4.2.2, it is apparent that the topics with the highest percentage of negative reviews, as determined by user ratings and sentiment analysis, are Sign Up and Sign In Problems, Network Connectivity, App Crashes, and App Configuration. Except for app configuration, the other three topics pose significant issues for users, as they prevent them from even opening the app.
Sign Up and Sign In Issues: Common problems in this category include system bugs that prevent users from registering or logging in even with correct credentials. Developers should address these bugs promptly and also provide seamless integration between web and mobile platforms to allow users to access their accounts across different devices. Enhancing user convenience by integrating sign-in options with Google, Facebook, or other social media accounts can also improve user experience and accessibility.
Network Connectivity: Users often face challenges with internet connectivity and server access, especially in areas with low bandwidth. Developers should optimize their apps to handle low-bandwidth situations effectively and consider server scalability to manage high traffic, particularly during peak times such as disasters. Implementing features that allow the app to function offline can be invaluable, especially since disasters can disrupt normal communication infrastructures.
App Crashes: Crashes are a significant barrier to app usability and can result from various factors, including poor memory management and incomplete testing. To address memory management issues, developers should optimize their code to efficiently use resources and regularly profile the app to identify and fix memory leaks. For testing, implementing a comprehensive testing protocol that includes unit, integration, and stress tests will help ensure that the app is robust and can handle real-world operating conditions without failure.
App Configuration: While users appreciate apps that offer customizable settings, frustrations arise when these configurations do not save correctly. To enhance user satisfaction, developers should ensure that the app reliably saves user settings and consider adding more flexible options to cater to individual preferences. Adding a help feature to guide users through setup and configuration processes can also improve the overall user experience.
Better Support Diverse App End Users and Platforms: A number of user feedback indicated poor support for diverse users of the disaster support apps. This included language aspects, as discussed above, as well as challenges with colors e.g.for colorblind users, size and complexity of screens, need to run apps on lower-spec devices e.g. more limited battery life, network communications, older handsets, etc. Disaster apps need to be better designed for use on a wide variety of handsets, operating system versions, low resource capabilities, low network bandwidth and capacity, and diverse end users of the app. They need careful testing with diverse users and platforms, including under diaster-induced situations e.g. stressed users, network interruptions, high loading of servers etc.
@@ 6.1 Internal Validity
App Filtering in Dataset: Constructing the dataset of disaster apps posed specific challenges that could threaten the validity of our study. The process of filtering risked excluding apps that contain disaster-related features, as they could be wrongly identified as false positives. To mitigate these threats, we implemented multiple verification layers by reading app descriptions, analyzing screenshots from Play Store and App Store, and directly downloading and evaluating the apps. Moreover, all three authors engaged in regular discussions to refine the selection of apps, ensuring that our dataset accurately represents relevant disaster management tools.
Cluster Naming: In our use of BERTopic for topic analysis, the algorithm automatically generates names for each cluster, which might not accurately reflect the core themes of the topics. To address this, we manually checked and, if necessary, renamed the clusters based on the evaluation of keywords and analysis of the k-closest reviews (where k=100 for this study) to their centroids. This manual intervention introduces a degree of subjectivity, which could affect the study’s validity. To counter this, we involved multiple reviewers in the process; three authors participated in reviews and discussions, and any disagreements were resolved using the negotiated agreement method [43], widely acknowledged in qualitative research.
@@ 6.2 Construct Validity
Fine-tuning the Models: The configuration and fine-tuning of BERTopic might have also influenced the results. The hyper-parameters such as the sentence embedding method, distance metric formula, and minimum topic size can significantly affect what the model identifies as a topic. Similarly, the configuration of the VADER model, particularly the sentiment thresholds set for classifying texts as positive (score >= 0.05), neutral (0.05 > score > -0.05), or negative (score <= -0.05), might impact the sentiment analysis. Since VADER is rule-based, there are no hyper-parameters to tune, but the thresholds chosen are crucial in determining the sentiment classification.
@@ 6.3 External Validity
Review Scraping: Data scraping was performed using both Play Store and App Store scrapers, employing batch scraping strategies to manage access limitations imposed by Google and Apple. We paused for 5 seconds after every 500 reviews, but despite these precautions, some reviews may still have been missed if access was denied due to scraping limits.
Translation: Due to limited resources, our translation efforts are focused solely on reviews from apps developed in non-English-speaking countries. While we acknowledge that apps from English-speaking regions can also provide reviews in various languages, we do not translate these in this research. This decision may impact the depth and breadth of our understanding of user feedback from different linguistic backgrounds.
@@ 6.4 Reliability
To enhance the reliability of our study and ensure reproducibility for future researchers, we have compiled a comprehensive replication package openly available on GitHub 999github.com/Syukrondzeko/natural-disaster-app-review-analysis, which includes the dataset used in our analysis and the source code of clustering and sentiment analysis models utilized. This repository is designed to facilitate replication of our study, providing step-by-step guidelines on how to use it, thus minimizing the risk of obtaining different results.
@@ 7 Conclusion
Disaster apps represent a group of apps that are critical for human survival. In this study we conducted a detailed analysis of 45 apps along with their reviews. The analysis of the app features resulted in creating a list of features allocated to four stages of disaster life cycle: preparation, response, recovery and mitigation. This pool of features will help emergency agencies and future app developers to get an overview of the features that can be included in disaster apps. The topics we have identified in RQ2 and the users reactions to these features from RQ3, will help future developers understand common issues that exist in disaster apps and take precautionary steps to prevent them in their apps. These RQ results also show which features were most preferred by users and by incorporating these features the future app developers can increase their user satisfaction. The recommendation we have provided can be applied for any crisis management app such as SOS apps, Cyber security apps and even health emergency apps.
8 Acknowledgement

### v2

@@ 5 Discussions and Recommendations
Expanding the Scope of Disaster Management Features in Mobile Apps: In disaster management applications, the focus often lies on functionalities that are active during or immediately after a disaster, such as real-time alerts, disaster maps, and updates. While these are undoubtedly crucial, there is a significant opportunity to expand the feature set to better prepare and support users throughout the entire disaster lifecycle. For example, early warning alerts, which are present in only 22% of the surveyed applications, can provide critical advance notice to users, allowing them to take necessary precautions and potentially save lives.
@@
Moreover, the recovery and mitigation phases are equally important, yet often underrepresented in current applications. Features like shelter information, safety confirmations, and recovery actions are vital for helping individuals locate assistance, confirm the safety of friends and family, and understand the next steps they can take. Enhancing the availability of these features could significantly improve the efficacy of disaster response and recovery efforts, addressing gaps that currently leave users under prepared or insufficiently supported after a disaster strikes. By integrating a more comprehensive set of features that span the pre- and post-disaster phases, developers can create more robust applications that cater to a wider range of user needs and scenarios.
@@
Internationalization: As shown in Table 2, among the various apps categorized based on their geographic coverage—22 global apps, 6 region/territorial apps, 20 national apps, and 8 local apps—only the global and regional apps consistently support the English language. Among the national apps from non-English speaking countries, such as ‘Anhaar’ from Mongolia, ‘Emergency Ready’ App from South Korea, ‘BD 999’ from Bangladesh, ‘Alertswiss’ from Switzerland, ‘NINA - Die Warn-App des BBK’ from Germany, and several from Japan including ‘NERV Disaster Prevention’, ‘Safety Tips’, ‘National Evacuation Center Guide’, and ‘Yurekuru Call’, only ‘Safety Tips’, ‘NERV Disaster Prevention’, and ‘Emergency Ready’ App offer English language support.
@@
These examples underscore the importance of incorporating at least one universally understood language, such as English, in disaster management apps. This feature is essential for ensuring the usability of the app especially for visitors and temporary residents. Providing quality translations or, alternatively, enabling features that assist with language translation, like text copying, can significantly enhance the functionality and user experience of these apps.
@@
Addressing Topics with the Highest Percentage of Negative Reviews: Based on user ratings and sentiment scores, the topics with the lowest values are Sign Up and Sign In Problems, Network Connectivity, App Crashes, Battery Drainage, and Watch Zone Functionality. Except for Battery Drainage and Watch Zone Functionality, the other three topics represent critical usability issues, as they prevent users from accessing or properly operating the app. These problems should be thoroughly resolved during the quality assurance (QA) testing stage to ensure a stable and functional user experience. Regarding Battery Drainage, developers could optimize background processes or implement efficient notification mechanisms to reduce power consumption. For Watch Zone Functionality, developers should allow users to customize or select their preferred watch zones, as this flexibility would improve the app’s relevance and usability across different geographic regions.
Better Support for Diverse App End Users and Platforms: Several user reviews highlighted inadequate support for diverse users of disaster apps. First, in terms of language, the app should provide a global language option, such as English, to accommodate travelers and foreign residents in different countries. Regarding accessibility, some users reported difficulties due to color vision impairments, as they were unable to distinguish certain colors on the map. In addition, small font sizes made it difficult for users with myopia to read the text. For audio alerts, users mentioned that excessively loud notifications disturbed individuals with sensitive hearing, while overly quiet alerts were ineffective for users with hearing impairments. Furthermore, some users reported compatibility issues with specific operating systems, preventing them from using the app. Lastly, many disaster apps require a stable or high-speed internet connection, which can limit access for users in rural or low-connectivity areas. Addressing these issues would make disaster apps more inclusive and accessible to a wider range of users.
@@ 6.1 Internal Validity
App Filtering in Dataset: Constructing the dataset of disaster apps posed specific challenges that could affect the validity of our study. The filtering process risked excluding relevant apps with disaster-related features due to misclassification. To address this issue, we downloaded and tested each app directly to verify its relevance and functionality. Furthermore, all three authors engaged in regular discussions to refine the inclusion criteria, ensuring that the final dataset accurately represents disaster management applications.
@@ 6.2 Construct Validity
Model Configuration: The configuration and fine-tuning of BERTopic might have also influenced the results. The hyper-parameters such as the sentence embedding method, distance metric formula, and minimum topic size can significantly affect what the model identifies as a topic. Similarly, the configuration of the VADER model, particularly the sentiment thresholds set for classifying texts as positive (score >= 0.05), neutral (0.05 > score > -0.05), or negative (score <= -0.05), might impact the sentiment analysis. Since VADER is rule-based, there are no hyper-parameters to tune, but the thresholds chosen are crucial in determining the sentiment classification.
@@ 6.3 External Validity
Review Scraping: Data scraping was performed using both the Google Play Store and the App Store scrapers, employing batch scraping strategies to manage access limitations imposed by Google and Apple. We paused for 5 seconds after every 500 reviews, but despite these precautions, some reviews may still have been missed if access was denied due to scraping limits.
Translation: Due to limited resources, our translation efforts focus solely on reviews from apps developed in non-English-speaking countries. While we acknowledge that apps from English-speaking regions may also contain reviews in various languages, we do not translate these in this study. This decision may affect the depth and breadth of our understanding of user feedback from different linguistic backgrounds, although the number of non-English reviews is relatively small.
@@ 6.4 Reliability
To enhance the reliability of our study and ensure reproducibility for future researchers, we have compiled a comprehensive replication package openly available on GitHub 101010github.com/Syukrondzeko/natural-disaster-app-review-analysis, which includes the dataset used in our analysis and the source code of clustering and sentiment analysis models utilized. This repository is designed to facilitate replication of our study, providing step-by-step guidelines on how to use it, thus minimizing the risk of obtaining different results.
@@ 7 Conclusion
Disaster apps represent a group of apps that are critical for human survival. In this study we conducted a detailed analysis of 44 apps along with their reviews. The analysis of the app features resulted in creating a list of features allocated to three stages of disaster lifecycle: preparation, response, and recovery. This pool of features will help emergency agencies and future app developers to get an overview of the features that can be included in disaster apps. The topics we have identified in RQ2 and the users reactions to these features from RQ3, will help future developers understand common issues that exist in disaster apps and take precautionary steps to prevent them in their apps. These RQ results also show which features were most preferred by users and by incorporating these features the future app developers can increase their user satisfaction.
8 Acknowledgment
@@
APPENDIX
Table A1: Region-Specific Disaster Apps
App Name Land Area Oceanic and Basin
USA
Canada
Europe
Atlantic
Ocean
Indian
Ocean
Pacific
Ocean
Central Pacific
Basin
Eastern Pacific
Basin
CodeRED Mobile Alert x x
FloodAlert Waterlevel Alerts x x
Global Storms x x x
Hurricane Tracker x x
SeaStorm Hurricane Tracker x x x
Tropical Hurricane Tracker x x x
Table A2: Overview of Disaster-Related Mobile Apps from App Store
No. App Name Downloads Rating Pricing Geographic Coverage Disaster Type
1 Alert2Me - Emergency Alerts 57 3.2 Free Global Multi-Disaster
2 Alert SA 129 2.3 Free Local Multi-Disaster
3 Alertswiss 522 3.7 Free National Multi-Disaster
4 Disaster Alert 26 3.8 Free Global Multi-Disaster
5 Earthquake 1.7K 4.8 Free with In-App Purchase Global Earthquake
6 Earthquake + Alerts Map & Info 790 4.6 Free with In-App Purchase Global Earthquake
7 Emergency Ready App 15 2.8 Free National Multi-Disaster
8 GeoNet 465 4.8 Free National Earthquake
9 Hazards Near Me NSW 1.1K+ 3.3 Free Local Multi-Disaster
10 LastQuake 104 4.7 Free Global Earthquake
11 My Earthquake Alerts 405 4.6 Free Global Earthquake
12 QuakeFeed Earthquake Tracker 1.1K 4.7 Free with In-App Purchase Global Earthquake
13 Safety tips 13 2.3 Free National Earthquake
14 SES Assistance QLD 6 1.7 Free National Flood
15 VIC Fires 6.4K 4.7 Paid National Fire
16 VicEmergency 614 2.9 Free National Multi-Disaster
Table A3: Overview of Disaster-Related Mobile Apps from Play Store
No. App Name Downloads Rating Pricing Geographic Coverage Disaster Type
1 Alert SA 100K+ 3.5 Free Local Multi-Disaster
2 Alert2Me - Emergency Alerts 10K+ 3.5 Free Global Multi-Disaster
3 Alertswiss 500K+ 4.1 Free National Multi-Disaster
4 Anhaar 10K+ 4.7 Free National Multi-Disaster
5 BD 999 10K+ 5.0 Free National Multi-Disaster
6 CodeRED Mobile Alert 500K+ 2.8 Free Regional Multi-Disaster
7 Disaster Alert 500K+ 4.1 Free Global Multi-Disaster
8 Earthquake 100K+ 4.1 Free with Ads Global Earthquake
9 Earthquake Alert! 1M+ 4.6 Free Global Earthquake
10 Earthquake + Alerts Map & Info 100K+ 4.1 Free with Ads and In-App Purchase Global Earthquake
11 Earthquakes Today 100K+ 4.4 Free with Ads and In-App Purchase Global Earthquake
12 Earthquakes Tracker 500K+ 4.5 Free with Ads and In-App Purchase Global Earthquake
13 Emergency 10K+ 4.2 Free National Multi-Disaster
14 Emergency Ready App 10K+ 3.5 Free National Multi-Disaster
15 Emergency: Severe Weather App 100K+ 3.7 Free National Multi-Disaster
16 FEMA 1M+ 3.8 Free National Multi-Disaster
17 FloodAlert Waterlevel Alerts 100K+ 4.2 Free with In-App Purchase Regional Flood
18 Floods Near Me NSW 50K+ 2.1 Free Local Flood
19 Global Storms 1M+ 4.4 Free with Ads and In-App Purchase Regional Wind-Related Disaster
20 Hazards Near Me NSW 1M+ 2.5 Free Local Multi-Disaster
21 Hazards Red Cross 50K+ 3.1 Free National Multi-Disaster
22 Hurricane & Typhoon Track 100K+ 3.8 Free with Ads Global Wind-Related Disaster
23 Hurricane Tracker 10K+ 4.3 Paid with In-App Purchase Regional Wind-Related Disaster
24 My Alert 50K+ 3.5 Free Global Multi-Disaster
25 My Earthquake Alerts 5M+ 4.7 Free with Ads Global Earthquake
26 My Hurricane Tracker 500K+ 4.7 Free with Ads Global Wind-Related Disaster
27 My Hurricane Tracker Pro 5K+ 4.7 Paid Global Wind-Related Disaster
28 NINA - Die Warn-App des BBK 5M+ 2.6 Free National Multi-Disaster
29 NERV Disaster Prevention 1M+ 4.6 Free with In-App Purchase National Multi-Disaster
30 Natural Disaster Monitor 50K+ 3.5 Free with Ads Global Multi-Disaster
31 National evacuation center guide 100K+ 3.3 Free National Multi-Disaster
32 SD Emergency 50K+ 3.5 Free Local Multi-Disaster
33 SES Assistance QLD 10K+ 3.1 Free Local Flood
34 Safety tips 100K+ 2.3 Free National Earthquake
35 SeaStorm Hurricane Tracker 10K+ 4.1 Free Regional Wind-Related Disaster
36 Tropical Hurricane Tracker 100K+ 4.9 Free with Ads and In-App Purchase Regional Wind-Related Disaster
37 VicEmergency 1M+ 3.3 Free Local Multi-Disaster
38 Volcanoes & Earthquakes 1M+ 4.4 Free with Ads and In-App Purchase Global Earthquake
39 Wind Map Hurricane Tracker 3D 1M+ 4.3 Free with Ads and In-App Purchase Global Wind-Related Disaster
40 Yurekuru Call 1M+ 4.4 Free with Ads and In-App Purchase National Earthquake